OpenStudy (anonymous):

The perimeter of a rectangle is 34 feet. Find the width of the rectangle if the length is 5 feet more than twice the width. Answer 4 feet 3 feet 5 feet 2 feet .

OpenStudy (anonymous):

Perimeter=2Length + 2Width Since we know that Length = 2Width + 5 we can work this out. I will refer to Length and Width as L and W respectively 34=2L+2W from the problem we know that L=2W+5 now we plug in what we just got for L 34=2(2W+5)+2W 34=4W+10+2W Isolate W 34-10=6W 24/6=W 4=W